There is a claim that on daily average, students spend 5.5 hours on social media with a standard deviation of 1.25. A researcher wants to test the claim at 5% level of significance.



1. Formulate the null and alternative hypothesis



2. Determine the test statistic to be used



3. Find the corresponding zvalue



4. Identify the rejection region




1
Expert's answer
2022-06-07T15:33:33-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=5.5"

"H_1:\mu\not=5.5"


2. This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


3. Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."


4. The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."


5. The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}"